### v1

# The coarse Baum-Connes conjecture with filtered coefficients and product metric spaces

Abstract.
Inspired by the quantitative $K$-theory, in this paper, we introduce the coarse Baum-Connes conjecture with filtered coefficients which generalizes the original conjecture. The are two advantages for the conjecture with filtered coefficients. Firstly, the routes toward the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients. Secondly, the class of metric spaces that satisfy the conjecture with filtered coefficients is closed under products and yet it is not known for the original conjecture. As an application, we show some new examples of product metric spaces for the coarse Baum-Connes conjecture.
The author was supported by NSFC12171156, 12271165, 12301154.
1. Introduction
The coarse Baum-Connes conjecture (cf. [8][19]) is dated back to the celebrated Atiyah-Singer index theorem which gives a formula by using topological index to compute the Fredholm index of an elliptic differential operator on a closed manifold. For a Riemannian manifold $M$ and an elliptic differential operator $D$ on $M$, Roe defined a higher index of $D$ which lives in the $K$-theory of a $C^{\ast}$-algebra, now called Roe algebra of $M$ (cf. [16]). Let us recall its definition. Choose a locally finite net $N_{M}$ in $M$, the Roe algebra of $M$ is defined to be the norm closure of the $\ast$-algebra consisting of all $(N_{M}\times N_{M})$-matrices $T$ with coefficients in $\mathfrak{K}(H)$ satisfying that $T_{x,y}=0$ when $d(x,y)>R$ for some $R>0$, where $\mathfrak{K}(H)$ is the algebra of compact operators on a Hilbert space $H$. The coarse Baum-Connes conjecture assert that the assembly map from the coarse $K$-homology of $M$ to the $K$-theory of Roe algebra of $M$ is an isomorphism (see Conjecture 3.23). Thus, the coarse Baum-Connes conjecture provides a topological algorithm for higher indexes. Moreover, the conjecture has important applications to the Novikov conjecture and the existence of Riemannian metric on a manifold with positive scalar curvature (cf. [9][18]).
There are three routes toward the coarse Baum-Connes conjecture. Roughly speaking, the first one is algebraic, in [21], Yu invented the quantitative $K$-theory to verify the conjecture holds for metric spaces with finite asymptotic dimension. The second one is geometric, in [22], Yu used the Dirac-dual-Dirac method to prove that the conjecture holds for metric spaces which admit a coarse embedding into Hilbert space. And the third one is topological, in [8], Higson and Roe using coarse homotopy and open cones to show the conjecture is true for non-positively curved metric spaces. On the other hand, there exist counterexamples to the coarse Baum-Connes conjecture (cf. [7][12]).
In [15], Oyono-Oyono and Yu systemized the quantitative $K$-theory for filtered $C^{\ast}$-algebras as defined below.
Definition 1.1.
([15, Definition 1.1]) Let $A$ be a $C^{\ast}$-algebra. A filtration of $A$ is a family of closed linear subspaces $(A_{r})_{r>0}$ such that
  • $A_{r}\subseteq A_{r^{\prime}}$, if $r\leq r^{\prime}$;
  • $A_{r}$ is closed by involution;
  • $A_{r}A_{r^{\prime}}\subseteq A_{r+r^{\prime}}$;
  • $\cup_{r>0}A_{r}$ is dense in $A$.
If $A$ is unital, we also require the unit of $A$ belongs to $A_{r}$ for any $r>0$. A $C^{\ast}$-algebra equipped with a filtration is called a filtered $C^{\ast}$-algebra.
As a refinement of the $K$-theory, the quantitative $K$-theory with order $r$ of $A$ considers homotopy equivalent classes of almost-projections or almost-unitaries in $A_{r}$.
Inspired by the quantitative $K$-theory, we define the Roe algebra of $M$ with filtered coefficients in $A$ to be the norm closure of the $\ast$-algebra consisting of all $(N_{M}\times N_{M})$-matrices $T$ with coefficients in $\mathfrak{K}(H)\otimes A_{r}$ satisfying that $T_{x,y}=0$ when $d(x,y)>R$ for some $r,R>0$ (see Definition 3.2). And we introduce the coarse Baum-Connes conjecture with filtered coefficients to compute the $K$-theory of Roe algebras with filtered coefficients by using of the $K$-theory of localization algebras with filtered coefficients (see Conjecture 3.23). When $A=\mathbb{C}$ with the trivial filtration $A_{r}=\mathbb{C}$, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is exactly the original conjecture. Moreover, the above-mentioned three ways to the original conjecture also work for the coarse Baum-Connes conjecture with filtered coefficients (see Section 5).
In order to explore the coarse Baum-Connes conjecture with filtered coefficients for product metric spaces, we introduce a notion of localization algebras along one direction with filtered coefficients as a bridge (see Definition 4.3). Then we obtain the following result (see also Theorem 4.19).
Theorem 1.2.
Let $X$, $Y$ be two proper metric spaces with bounded geometry and $A$ be a filtered $C^{\ast}$-algebra. If the following two conditions are satisfied:
  1. (1)
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(Y,A)$ holds for $X$;
  2. (2)
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $Y$, where $\square$ is the separated coarse union of a sequence of single points.
Then the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X\times Y$.
In particular, if the coarse Baum-Connes conjecture with filtered coefficients holds for $X$ and $Y$, then the coarse Baum-Connes conjecture with filtered coefficients holds for $X\times Y$.
The above theorem is a coarse analogue of H. Oyono-Oyono’s result in [14] stating that the Baum-Connes conjecture with coefficients is closed under the direct product of groups.
In [5], Fukaya and Oguni proved that the coarse Baum-Connes conjecture holds for product metric spaces of certain groups by studying the coronae of products. In generally, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture, it is not known whether the conjecture is true or not for $X\times Y$. Take $A=\mathbb{C}$ in Theorem 1.2, we have the following corollary (see also Corollary 4.20) which not only recovers Fukaya and Oguni’s result but also provides new examples of products satisfying the conjecture (see Section 5).
Corollary 1.3.
Let $X$ and $Y$ be two proper metric spaces with bounded geometry. If the followings are satisfied:
  1. (1)
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}(Y)$;
  2. (2)
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$.
Then the coarse Baum-Connes conjecture holds for $X\times Y$.
In particular, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture with filtered coefficients, then $X\times Y$ satisfies the coarse Baum-Connes conjecture.
Besides, the coarse Baum-Connes conjecture with filtered coefficients is also applied to the quantitative coarse Baum-Connes conjecture which is a refinement of the original conjecture, we will explore this connection in the subsequent paper (cf. [23]).
The paper is organized as follows. In Section 2, we recall some ingredients from the $K$-theory of $C^{\ast}$-algebras. In Section 3, we introduce Roe algebras with filtered coefficients, localization algebras with filtered coefficients and discuss their properties, finally, we formulate the coarse Baum-Connes conjecture with filtered coefficients. In Section 4, we give the coarse Baum-Connes conjecture along one direction for product metric spaces, then build a connection between it and the conjecture with filtered coefficients, finally, we prove Theorem 1.2. In Section 5, we apply our results to a number of metric spaces .
2. Preliminaries
In this section, we recall three important techniques for computing the $K$-theory of $C^{\ast}$-algebras. The main reference is Willett and Yu’s book [18].
Lemma 2.1.
[18, Proposition 2.7.5] Let $\alpha:C\rightarrow D$ be a $\ast$-homomorphism and $v$ be a partial isometry in the multiplier algebra of $D$ satisfying that $\alpha(c)v^{\ast}v=\alpha(c)$ for all $c\in C$. Define $\beta:C\rightarrow D$ by $\beta(c)=v\alpha(c)v^{\ast}$. Then $\alpha$ and $\beta$ induce the same morphism on $K$-theory.
Definition 2.2.
[18, Definition 2.7.11] A $C^{\ast}$-algebra $B$ is called to be quasi-stable if for all positive integer $n$ there exists an isometry $v$ in the multiplier algebra of $M_{n}(A)$ such that $vv^{\ast}$ is the matrix unit $e_{11}$.
One of the advantages of quasi-stable $C^{\ast}$-algebras is as follow.
Lemma 2.3.
[18, Lemma 12.4.3] Let $B$ be a quasi-stable $C^{\ast}$-algebra and let $C_{ub}([0,\infty),B)$ be the $C^{\ast}$-algebra consisting of all bounded, uniformly continuous functions from $[0,\infty)$ to $B$, then the evaluation at zero map
$e:C_{ub}([0,\infty),B)\rightarrow B$
induces an isomorphism on $K$-theory.
The following Mayer-Vietoris six-term exact sequence is a key technique to compute the $K$-theory of $C^{\ast}$-algebras by their ideals, for the proof, please refer to [18, Proposition 2.7.15] or [10].
Lemma 2.4.
Let $I_{1}$ and $I_{2}$ be two closed ideals in a $C^{\ast}$-algebra $B$. Assume $I_{1}+I_{2}$ is dense in $B$. Then there exists a Mayer-Vietoris six-term exact sequence:
${K_{0}(I_{1}\cap I_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(I_{1})\oplus K_{0}(I_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(B)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(I_{1})\oplus K_{1}(I_{2})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(I_{1}\cap I_{2}),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
which is natural for triples $(I_{1},I_{2},B)$.
Now, Let us fix some notations in the rest of this paper. Let $(X,d)$ be a proper metric space (the properness means every closed ball in $X$ is compact). Since that proper metric spaces are separable, thus we can choose a countable dense subset $Z_{X}$ in $X$. And let $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. Fixed $H$ as a separable Hilbert space.
3. The coarse Baum-Connes conjecture with filtered coefficients
In this section, we will introduce the coarse Baum-Connes conjecture with filtered coefficients which generalizes the original conjecture.
3.1. Roe algebras with filtered coefficients
The Roe algebra was introduced by Roe in [17] motivated by the index theory on open manifolds [16]. In this subsection, we introduce a notion of Roe algebra with filtered coefficients in a filtered $C^{\ast}$-algebra.
Definition 3.1.
Let $X$, $Y$ be two proper metric spaces and $Z_{X}$, $Z_{Y}$ be two countable dense subsets of $X$, $Y$, respectively. Let $\chi$ be the characteristic function.
  1. (1)
    For a bounded operator $T:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$, the support of $T$, denoted by $\operatorname{supp}(T)$, is defined to be the set consisting of all $(x,y)\in X\times Y$ satisfying that $(\chi_{V}\otimes I\otimes I)T(\chi_{U}\otimes I\otimes I)\neq 0$ for any open neighborhoods $U$ and $V$ of $x$ and $y$, respectively.
  2. (2)
    For a bounded operator $T$ on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$, the propagation of $T$, denoted by $\operatorname{prop}(T)$, is defined to be $\sup\{d(x,y):(x,y)\in\operatorname{supp}(T)\}$.
Definition 3.2.
The Roe algebra of $X$ with filtered coefficients in $A$, denoted by $C^{\ast}_{f}(X,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all operators $T$ on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$ satisfying that
  • $T$ has finite propagation;
  • there exists $s>0$ such that for any compact subset $K\subseteq X$, the operators $(\chi_{K}\otimes I\otimes I)T$ and $T(\chi_{K}\otimes I\otimes I)$ are in $\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$, where $\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)$ is the algebra of compact operators on $\ell^{2}(Z_{X})\otimes H$.
Remark 3.3.
If $A$ is a $C^{\ast}$-algebra is equipped with the trivial filtration, namely, $A_{r}=A$ for any $r>0$, then $C_{f}^{\ast}(X,A)$ is the Roe algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}(X,A)$. In particular, when $A=\mathbb{C}$, the $C^{\ast}$-algebra $C_{f}^{\ast}(X,\mathbb{C})$ or $C^{\ast}(X)$ for short is so called the Roe algebra of $X$ (cf. [17]).
Now, we discuss the functoriality on proper metric spaces for Roe algebras with filtered coefficients.
Definition 3.4.
Let $X$ and $Y$ be two proper metric spaces. A map $g:X\rightarrow Y$ is called a coarse map, if
  • the inverse image of $g$ of any compact subset is pre-compact;
  • for any $R>0$, there exists $S>0$ such that $\sup\{d(g(x),g(y)):d(x,y)\leq R\}\leq S$.
For a coarse map $g$, define a function $w_{g}:\mathbb{R}^{+}\rightarrow\mathbb{R}^{+}$ by
$w_{g}(R)=\sup\{d(g(x),g(y)):\>\>d(x,y)\leq R\}.$
Definition 3.5.
Two proper metric spaces $X$ and $Y$ are called to be coarsely equivalent, if there exists two coarse maps $g:X\rightarrow Y$, $h:Y\rightarrow X$ and a constant $c>0$ such that
$\max\{d(hg(x),x),d(gh(y),y)\}\leq c,$
for any $x\in X$ and $y\in Y$.
For a coarse map $g:X\rightarrow Y$ and any $\delta>0$. Choose a disjoint Borel cover $(U_{i})_{i\in\mathbb{N}}$ of $Y$ with diameter less than $\delta$ and every $U_{i}$ has non-empty interior, then there exists a family of isometries $V_{i}:\ell^{2}(g^{-1}(U_{i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{i}\cap Z_{Y})\otimes H$. Let
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
Then $V_{g}$ is an isometry and
(3.1) $\operatorname{supp}(V_{g})\subseteq\{(x,y)\in X\times Y:\>d(g(x),y)<\delta\}.$
Thus we have the following lemma.
Lemma 3.6.
For a coarse map $g:X\rightarrow Y$ and $\delta>0$, the above isometry $V_{g}$ induces a $\ast$-homomorphism
$ad_{g}:C_{f}^{\ast}(X,A)\rightarrow C_{f}^{\ast}(Y,A),$
defined by
$ad_{g}(T)=V_{g}TV^{\ast}_{g},$
that satisfies
$\operatorname{prop}(ad_{g}(T))\leq w_{g}(\operatorname{prop}(T))+2\delta.$
Moreover, another $\ast$-homomorphism $ad^{\prime}_{g}$ defined by another isometry $V^{\prime}_{g}$ that satisfies the condition (3.1) induces the same homomorphism as $ad_{g}$ on $K$-theory.
Proof.
Firstly, if $(y,y^{\prime})\in\operatorname{supp}(V_{g}TV^{\ast}_{g})$, then there exist $x,x^{\prime}\in X$ such that $(y,x)\in\operatorname{supp}(V^{\ast}_{g})$, $(x,x^{\prime})\in\operatorname{supp}(T)$ and $(x^{\prime},y^{\prime})\in\operatorname{supp}(V_{g})$. Thus
$d(y,y^{\prime})\leq d(y,g(x))+d(g(x),g(x^{\prime}))+d(g(x^{\prime}),y^{\prime})\leq w_{g}(\operatorname{prop}(T))+2\delta,$
which implies $\operatorname{prop}(ad_{g}(T))\leq w_{g}(\operatorname{prop}(T))+2\delta$.
Secondly, for any compact subset $K$ of $Y$, let
$K^{\prime}=\{x\in X:\>\>\textrm{there exists $y\in K$ such that $(y,x)\in\operatorname{supp}(V^{\ast}_{g})$}\}.$
Then we have $K^{\prime}\subseteq g^{-1}(\overline{B_{\delta}(K)})$ since $d(g(x),y)<\delta$ for any $(y,x)\in\operatorname{supp}(V^{\ast}_{g})$, where $\overline{B_{\delta}(K)}=\{y\in Y:d(y,K)\leq\delta\}$ which is compact by the properness of $Y$. Thus $K^{\prime}$ is a compact subset of $X$ by the properness of $g$. Moreover, we have that $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)=V_{g}T(\chi_{K^{\prime}}\otimes I\otimes I)V^{\ast}_{g}(\chi_{K}\otimes I\otimes I)$ which implies $V_{g}TV^{\ast}_{g}(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$ since $T(\chi_{K^{\prime}}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{s}$. Similarly, we can obtain the similar result for $(\chi_{K}\otimes I\otimes I)V_{g}TV^{\ast}_{g}$. In conclusion, $ad_{g}$ is well-defined.
Finally, if there exists another isometry $V^{\prime}_{g}:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ satisfying the condition (3.1). Then by similar arguments as above, $V_{g}V^{\ast}_{g}$, $V_{g}^{\prime}V^{\prime\ast}_{g}$, $V^{\prime}_{g}V^{\ast}_{g}$ are all multipliers of $C_{f}^{\ast}(Y,A)$. And
$\begin{pmatrix}0&0\\ 0&V^{\prime}_{g}TV^{\prime\ast}_{g}\end{pmatrix}=\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix}\begin{pmatrix}V_{g}TV^{\ast}_{g}&0\\ 0&0\end{pmatrix}\begin{pmatrix}I-V_{g}V^{\ast}_{g}&V_{g}V^{\prime\ast}_{g}\\ V^{\prime}_{g}V^{\ast}_{g}&I-V^{\prime}_{g}V^{\prime\ast}_{g}\end{pmatrix}$
which implies that $ad^{\prime}_{g}$ and $ad_{g}$ induce the same homomorphism on $K$-theory by Lemma 2.1 because their images are unitarily equivalent by a unitary multiplier. ∎
If $g$ is the identity map on $X$, then we can choose $V_{g}$ to be a unitary, thus we obtain the following corollary.
Corollary 3.7.
For different choices of the dense subset $Z_{X}$ of $X$, Roe algebras with filtered coefficients are non-canonically isomorphic and their $K$-theory are canonically isomorphic.
Corollary 3.8.
If two proper metric spaces $X$ and $Y$ are coarsely equivalent, then
$C_{f}^{\ast}(X,A)\cong C_{f}^{\ast}(Y,A)$
for any filtered $C^{\ast}$-algebras $A$.
Proof.
By the definition of coarse equivalence (see Definition 3.5), there exist two coarse maps $g:X\rightarrow Y$, $h:Y\rightarrow X$ and a constant $c>0$ such that
$\max\{d(hg(x),x),d(gh(y),y)\}\leq c.$
Thus we can choose a disjoint Borel cover $(U_{i})_{i\in\mathbb{N}}$ of $Y$ with every $U_{i}$ has non-empty interior such that $g^{-1}(U_{i})\neq\emptyset$ and the diameter of $U_{i}$ is less than some positive number $\delta$ for any $i\in\mathbb{N}$. Let $Z_{Y}$ be a countable dense subset of $Y$ and $Z_{X}$ be a countable dense subset of $X$ with $Z_{X}\cap g^{-1}(U_{i})\neq\emptyset$ for all $i\in\mathbb{N}$. Then for every $i\in\mathbb{N}$, there exists a family of unitaries $V_{i}:\ell^{2}(g^{-1}(U_{i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{i}\cap Z_{Y})\otimes H$. Define
$V_{g}=\bigoplus_{i\in\mathbb{N}}V_{i}\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A}.$
Then $V_{g}$ is a unitary and we have that
$\operatorname{supp}(V_{g})\subseteq\{(x,y)\in X\times Y:\>d(g(x),y)<\delta\}.$
Thus by Lemma 3.6, $ad_{g}$ is an isomorphism from $C_{f}^{\ast}(X,A)$ to $C_{f}^{\ast}(Y,A)$. ∎
In the end of this subsection, we discuss two properties of Roe algebras with filtered coefficients. Firstly, we show that they are filtered $C^{\ast}$-algebas equipped with the following filtrations.
Lemma 3.9.
The algebra $C_{f}^{\ast}(X,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration $C_{f}^{\ast}(X,A)_{r}$ consisting of all operators $T\in C_{f}^{\ast}(X,A)$ such that $\operatorname{prop}(T)\leq r$ and $(\chi_{K}\otimes I\otimes I)T,T(\chi_{K}\otimes I\otimes I)\in\mathfrak{K}(\ell^{2}(Z_{X})\otimes H)\otimes A_{r}$ for any compact subset $K\subseteq X$.
Proof.
For any two operators $T,S\in C_{f}^{\ast}(X,A)$, we have that $\operatorname{supp}(TS)$ is contained in the closure of the set
$\{(x,y)\in X\times X:\>\>\textrm{there exists $z\in X$ such that $(x,z)\in\operatorname{supp}(S)$, $(z,y)\in\operatorname{supp}(T)$}\}$
which implies
$\operatorname{prop}(TS)\leq\operatorname{prop}(T)+\operatorname{prop}(S).$
Besides, since that $(A_{r})_{r>0}$ is a filtration of $A$, thus $(C_{f}^{\ast}(X,A)_{r})_{r>0}$ is a filtration of $C_{f}^{\ast}(X,A)$. ∎
Secondly, we show that Roe algebras with filtered coefficients are quasi-stable.
Lemma 3.10.
Let $X$ be a proper metric space and $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. Then the Roe algebra $C_{f}^{\ast}(X,A)$ is quasi-stable.
Proof.
Let $Z_{X}$ be a countable dense subset of $X$ and $H$ be a separable Hilbert space. For any positive integer $n$, define an isometry $V:H^{n}\rightarrow H\oplus H^{n-1}$ with the first coordinate $H$ as its image. Then $I\otimes V\otimes I:\ell^{2}(Z_{X})\otimes H^{n}\otimes H_{A}\rightarrow\ell^{2}(Z_{X})\otimes H^{n}\otimes H_{A}$ is an isometry in the multiplier algebra of $M_{n}(C_{f}^{\ast}(X,A))$ with the propagation zero. Moreover, we have that $(I\otimes V\otimes I)(I\otimes V\otimes I)^{\ast}=I\otimes VV^{\ast}\otimes I=e_{11}$. Thus $C_{f}^{\ast}(X,A)$ is quasi-stable. ∎
3.2. Localization algebras with filtered coefficients
Yu in [20] defined a notion of the localization algebra whose $K$-theory provides a model for $K$-homology. In this subsection, we consider the localization algebra with filtered coefficients.
Definition 3.11.
The localization algebra of $X$ with filtered coefficients in $A$, denoted by $C^{\ast}_{L,f}(X,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C^{\ast}_{f}(X,A)$ satisfying that
$\lim_{t\rightarrow\infty}\operatorname{prop}(u(t))=0.$
Remark 3.12.
If a $C^{\ast}$-algebra $A$ is equipped with the trivial filtration, then $C^{\ast}_{L,f}(X,A)$ is the localization algebra of $X$ with coefficients in $A$, also denoted by $C^{\ast}_{L}(X,A)$. In particular, when $A=\mathbb{C}$, the $C^{\ast}$-algebra $C^{\ast}_{L,f}(X,\mathbb{C})$ or $C^{\ast}_{L}(X)$ for short is so called the localization algebra of $X$ (cf. [20]).
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a deceasing sequence of positive numbers with limit zero. Choose a sequence of Borel covers $\{\mathcal{U}_{k}\}_{k\in\mathbb{N}}$ of $Y$ such that
  • every cover $\mathcal{U}_{k}$ is disjoint, i.e. $U_{k,i}\cap U_{k,j}=\emptyset$ for different $i,j$;
  • every $U_{k,i}\in\mathcal{U}_{k}$ has non-empty interior;
  • the diameter of $\mathcal{U}_{k}$ is less than $\delta_{k}$ for any $k\in\mathbb{N}$;
  • $\mathcal{U}_{k+1}$ refines $\mathcal{U}_{k}$, i.e. every set in $\mathcal{U}_{k+1}$ is contained in some set in $\mathcal{U}_{k}$ for all $k\in\mathbb{N}$.
Then $\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ is an infinitely dimensional separable Hilbert space for every set $U_{k,i}\in\mathcal{U}_{k}$. Thus we can define an isometry $V_{k,i}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ with infinitely dimensional cokernel. By the fourth condition as above, there exists $\{U_{k+1,ij}\}_{1\leq j\leq n}\subseteq\mathcal{U}_{k+1}$ such that $\cup^{n}_{j=1}U_{k+1,ij}=U_{k,i}$. Then we can choose $n$ number of isometries $V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k+1,ij})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k+1,ij}\cap Z_{Y})\otimes H$ with infinitely dimensional cokernel. Let $V_{k+1,i}=\bigoplus^{n}_{j=1}V_{k+1,ij}:\ell^{2}(g^{-1}(U_{k,i})\cap Z_{X})\otimes H\rightarrow\ell^{2}(U_{k,i}\cap Z_{Y})\otimes H$ that is an isometry with infinitely dimensional cokernel. Let $W_{k,i}$ be a unitary from the cokernel of $V_{k,i}$ to the cokernel of $V_{k+1,i}$ and let $W^{\prime}_{k,i}=V_{k+1,i}V^{\ast}_{k,i}+0\oplus W_{k,i}$ which is a unitary on $\ell^{2}(U_{k,i}\cap Z_{Y})$. Let $h:S^{1}\rightarrow[0,1)$ be the Borel inverse function of the continuous function $t^{\prime}\mapsto e^{2\pi it^{\prime}}$. Define $\gamma_{k,i}$ to be a path of isometries by $\gamma_{k,i}(t)=e^{2\pi ith(W^{\prime}_{k,i})}V_{k,i}$. Then $\gamma_{k,i}$ is a $2\pi$-Lipschitz path connecting $V_{k,i}$ with $V_{k+1,i}$ and $\operatorname{supp}(\gamma_{k,i}(t))\subseteq g^{-1}(U_{k,i})\times U_{k,i}$. Define
$\gamma_{k}(t)=\bigoplus_{i}\gamma_{k,i}(t):\ell^{2}(Z_{X})\otimes H\rightarrow\ell^{2}(Z_{Y})\otimes H$
for $t\in[0,1]$. Then for $s\in[0,\infty)$, define
$V_{g}(s)=\gamma_{k}(s-k)\otimes I:\ell^{2}(Z_{X})\otimes H\otimes H_{A}\rightarrow\ell^{2}(Z_{Y})\otimes H\otimes H_{A},\>\>\textrm{if $s\in[k,k+1]$}.$
Then $V_{g}$ is a bounded and uniformly continuous function satisfying that
(3.2) $\operatorname{supp}(V_{g}(s))\subseteq\{(x,y)\in X\times Y:d(g(x),y)<\delta_{k}\},\>\>\textrm{if $s\in[k,k+1]$}.$
Thus by Lemma 3.6, we have the following lemma.
Lemma 3.13.
Let $g:X\rightarrow Y$ be a uniformly continuous coarse map and $\{\delta_{k}\}_{k\in\mathbb{N}}$ be a deceasing sequence of positive numbers with limit zero, then $V_{g}$ as defined above induces a $\ast$-homomorphism
$Ad_{g}:C^{\ast}_{L,f}(X,A)\rightarrow C^{\ast}_{L,f}(Y,A)$
defined by
$Ad_{g}(u)(t)=V_{g}(t)u(t)V^{\ast}_{g}(t),$
that satisfies
$\operatorname{prop}(Ad_{g}(u)(t))\leq w_{g}(\operatorname{prop}(u(t)))+2\delta_{k}\>\>\textrm{for $t\in[k,k+1]$}.$
Moreover, for another bounded and uniformly continuous function $V^{\prime}_{g}$ which maps $[0,\infty)$ to the set of all isometries on $\ell^{2}(Z_{X})\otimes H\otimes H_{A}$ and satisfies the condition (3.2), then $Ad^{\prime}_{g}$ defined by $V^{\prime}_{g}$ induces the same homomorphism as $Ad_{g}$ on $K$-theory.
If replacing the Hilbert space $H$ by $H^{\infty}=\oplus_{i\geq 0}H$ in the definition of the localization algebra with filtered coefficients, then we obtain a $C^{\ast}$-algebra, denoted by $C^{\ast}_{L,f}(X,A,H^{\infty})$. Thus by Lemma 2.1, we have the following lemma.
Lemma 3.14.
Define $V:H\rightarrow H^{\infty}$ by $V(\xi)=(\xi,0,\cdots,0)$, then the map from $C^{\ast}_{L,f}(X,A)$ to $C^{\ast}_{L,f}(X,A,H^{\infty})$ defined by $u(t)\mapsto(I\otimes V\otimes I)u(t)(I\otimes V^{\ast}\otimes I)$ induces an isomorphism on the $K$-theory.
Next, we will show some homological properties of the $K$-theory of the localization algebras with filtered coefficients.
Definition 3.15.
[20, Definition 3.5] Let $X$, $Y$ be two proper metric spaces and $g$, $h$ be two Lipschitz maps from $X$ to $Y$. A continuous homotopy $F(t,x)$ $(t\in[0,1])$ between $g$ and $h$ is said to be strongly Lipschitz if
  1. (1)
    $F(t,x)$ is a coarse map from $X$ to $Y$ for each $t\in[0,1]$;
  2. (2)
    $d(F(t,x),F(t,y))\leq Cd(x,y)$ for all $x,y\in X$ and $t\in[0,1]$, where $C$ is a constant (called Lipschitz constant of $F$);
  3. (3)
    for any $\varepsilon>0$, there exists $\delta>0$ such that $d(F(t_{1},x),F(t_{2},x))<\varepsilon$ for all $x\in X$ if $\|t_{1}-t_{2}\|<\delta$;
  4. (4)
    $F(0,x)=g(x)$ and $F(1,x)=h(x)$ for all $x\in X$.
Definition 3.16.
[20, Definition 3.6] $X$ is said to be strongly Lipschitz homotopy equivalent to $Y$ if there exist two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively.
The $K$-theory of localization algebras with filtered coefficients is computable because of the following two lemmas, the proofs are coming from Yu’s paper [20].
Lemma 3.17.
If $X$ is strongly Lipschitz homotopy equivalent to $Y$, then $K_{\ast}(C^{\ast}_{L,f}(X,A))$ is naturally isomorphic to $K_{\ast}(C^{\ast}_{L,f}(Y,A))$.
Proof.
By the assumption, there exists two Lipschitz maps $g:X\rightarrow Y$ and $h:Y\rightarrow X$ such that $hg$ and $gh$ are strongly Lipschitz homotopic to $id_{X}$ and $id_{Y}$, respectively. It is sufficient to prove that $Ad_{hg}$ induces the identity map on $K_{\ast}(C^{\ast}_{L,f}(X,A))$ and similar to $Ad_{gh}$. Let $F(t,x)$ be a strongly Lipschitz homotopy such that $F(0,\cdot)=hg$ and $F(1,\cdot)=id_{X}$. Then we can choose a sequence of non-negative numbers $\{t_{i,j}\}_{i,j\geq 0}$ in $[0,1]$ such that
  • $t_{0,j}=0$ and $t_{i,j}\leq t_{i+1,j}$;
  • for each $j$, there exists $N_{j}\geq 0$ such that $t_{i,j}=1$ for all $i\geq N_{j}$;
  • $d(F(t_{i,j},x),F(t_{i+1,j},x))\leq 1/(j+1)$ and $d(F(t_{i,j},x),F(t_{i,j+1},x))\leq 1/(j+1)$ for all $x\in X$.
By the construction before Lemma 3.13, there exists a sequence of multipliers $V_{F(t_{i,j},\cdot)}$ of $C^{\ast}_{L,f}(X,A)$ such that
$\operatorname{supp}(V_{F(t_{i,j},\cdot)}(t))\subseteq\{(x,x^{\prime})\in X\times X:d(F(t_{i,j},x),x^{\prime})<1/(1+i+j)\}$
for any $i,j\geq 0$. Define
$V_{i}(t)=R(t-j)(V_{F(t_{i,j},\cdot)}(t)\oplus V_{F(t_{i,j+1},\cdot)}(t))R^{\ast}(t-j),\>t\in[j,j+1],$
where
$R(t)=\begin{pmatrix}\cos(\pi t/2)&\sin(\pi t/2)\\ -\sin(\pi t/2)&\cos(\pi t/2)\end{pmatrix}.$
Now we prove the lemma for =1$\ast$=1∗ = 1 and by a suspension argument, we can similarly prove for =0$\ast$=0∗ = 0. Let $[u]\in K_{1}(C^{\ast}_{L,f}(X,A))$, let us assume $u\in C^{\ast}_{L,f}(X,A)^{+}$ and let $W=u\oplus I$. Define the following three unitaries:
$\begin{split}&a_{1}=\bigoplus_{i\geq 0}(V_{i}(t)((u(t)-I)\oplus 0)V^{\ast}_{i}(t)+I)\cdot W^{\ast};\\ &a_{2}=\bigoplus_{i\geq 0}(V_{i+1}(t)((u(t)-I)\oplus 0)V^{\ast}_{i+1}(t)+I)\cdot W^{\ast};\\ &a_{3}=I\bigoplus_{i\geq 1}(V_{i}(t)((u(t)-I)\oplus 0)V^{\ast}_{i}(t)+I)\cdot W^{\ast}.\end{split}$
Then $a_{i}\in M_{2}(C^{\ast}_{L,f}(X,A,H^{\infty}))^{+}$ since that $\operatorname{prop}(a_{i})\leq(C+1)\operatorname{prop}(u(t))+1/(j+1)$ for $t\in[j,j+1]$ and $V_{i}(t)$ acts on the subspace $H_{A}$ as the identity operator, where $C$ is Lipschitz constant of $F$. Moreover, since $d(F(t_{i,j},x),F(t_{i+1,j},x))\leq 1/(j+1)$, thus $[a_{1}]=[a_{2}]$ in $K_{1}(C^{\ast}_{L,f}(X,A))$ (the proof is similar to the last section of the proof of Lemma 3.6). Let $V:H^{\infty}\rightarrow H^{\infty},\>(h_{1},h_{2},\cdots)\mapsto(0,h_{1},\cdots)$, then $I_{\ell^{2}(Z_{X})}\otimes V\otimes I_{H_{A}}$ is an isometric multiplier of $C^{\ast}_{L,f}(X,A,H^{\infty})$. And
$a_{3}=(I_{\ell^{2}(Z_{X})}\otimes V\otimes I_{H_{A}})(a_{2}-I)(I_{\ell^{2}(Z_{X})}\otimes V^{\ast}\otimes I_{H_{A}})+I,$
thus $[a_{2}]=[a_{3}]$ in $K_{1}(C^{\ast}_{L,f}(X,A))$ by Lemma 2.1. Then, we have that
$[a_{1}a^{-1}_{3}]=[(V_{i}(t)((u(t)-I)\oplus 0)V^{\ast}_{i}(t)+I)\cdot W^{\ast}\bigoplus_{i\geq 1}I]=[Ad_{hg}(u)W^{\ast}\bigoplus_{i\geq 1}I]$
which is equal to $[I]$. Thus $Ad_{hg,\ast}=id$ on $K_{1}(C^{\ast}_{L,f}(X,a))$ by Lemma 3.14. ∎
Definition 3.18.
A cover $\{X_{1},X_{2}\}$ of $X$ is called to be uniformly excisive, if for any $\varepsilon_{t}\geq 0$ with $\lim_{t\rightarrow\infty}\varepsilon_{t}=0$, there exist $\delta_{t}\geq 0$ with $\lim_{t\rightarrow\infty}\delta_{t}=0$ such that
$B_{\varepsilon_{t}}(X_{1})\cap B_{\varepsilon_{t}}(X_{2})\subseteq B_{\delta_{t}}(X_{1}\cap X_{2}),$
for any $t\geq 0$, where $B_{\varepsilon_{t}}(X_{i})$ is the $\varepsilon_{t}$-neighborhood of $X_{i}$.
Lemma 3.19.
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2},X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2},X_{1}\cap X_{2}$, respectively, then there is a Mayer-Vietoris six-term exact sequence
${K_{0}(C^{\ast}_{L,f}(Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(C^{\ast}_{L,f}(X_{1},A))\oplus K_{0}(C^{\ast}_{L,f}(X_{2},A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(C^{\ast}_{L,f}(X,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,f}(X,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,f}(X_{1},A))\oplus K_{1}(C^{\ast}_{L,f}(X_{2},A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,f}(Y,A)),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where $Y=X_{1}\cap X_{2}$.
Proof.
For $i=1,2$, let $C^{\ast}_{L,f}(X,A;X_{i})$ be a closed subalgebra of $C^{\ast}_{L,f}(X,A)$ generated by all elements $u$ satisfying that there exists $c_{t}\geq 0$ with $\lim_{t\rightarrow\infty}c_{t}=0$ such that
$\operatorname{supp}(u(t))\subseteq\{(x,x^{\prime})\in X\times X:d((x,x^{\prime}),X_{i}\times X_{i})\leq c_{t}\},$
for all $t\in[0,\infty)$. Then there exists a natural inclusion map
$\tau:C^{\ast}_{L,f}(X_{i},A)\rightarrow C^{\ast}_{L,f}(X,A,X_{i}).$
Now we prove that $\tau$ induces an isomorphism on the $K$-theory. By the assumption, there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhood $B_{\varepsilon_{0}}(X_{i})$ of $X_{i}$ is strongly Lipschitz homotopy equivalent to $X_{i}$. For any element $u\in C^{\ast}_{L,f}(X,A;X_{i})$, there exists $t_{0}>0$ such that
$\operatorname{supp}(u(t+t_{0}))\subseteq B_{\varepsilon_{0}}(X_{i})\times B_{\varepsilon_{0}}(X_{i}),$
for all $t\geq 0$. Let $v(t)=u(t+t_{0})$, then $u$ and $v$ are homotopic equivalent in $C^{\ast}_{L,f}(X,A;X_{i})$ since that $u$ is uniformly continuous. Thus the inclusion map from $C^{\ast}_{L,f}(B_{\varepsilon_{0}}(X_{i}),A)$ to $C^{\ast}_{L,f}(X,A;X_{i})$ induces an isomorphism on the $K$-theory. Then $\tau_{\ast}:K_{\ast}(C^{\ast}_{L,f}(X_{i},A))\rightarrow K_{\ast}(C^{\ast}_{L,f}(X,A;X_{i}))$ is an isomorphism by Lemma 3.17.
On the other hand, $C^{\ast}_{L,f}(X,A;X_{i})$ is a ideal of $C^{\ast}_{L,f}(X,A)$ for $i=1,2$. Moreover, we have $C^{\ast}_{L,f}(X,A;X_{1})+C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A)$ and $C^{\ast}_{L,f}(X,A;X_{1})\cap C^{\ast}_{L,f}(X,A;X_{2})=C^{\ast}_{L,f}(X,A;X_{1}\cap X_{2})$. Thus we completed the proof by Lemma 2.4. ∎
3.3. The coarse Baum-Connes conjecture with filtered coefficients
Before building the coarse Baum-Connes conjecture with filtered coefficients, we need introduce a coarsening space for a metric space by Rips complexes.
Let $(N,d_{N})$ be a locally finite metric space, namely, any ball in $N$ just contains finite elements. For any $k>0$, the Rips complex of $N$ at scale $k$, denoted by $P_{k}(N)$, is a simplicial complex whose set of vertices is $N$ and a subset $\{z_{0},\cdots,z_{n}\}\subseteq N$ spans an $n$-simplex in paper $P_{k}(N)$ if and only if $d_{N}(z_{i},z_{j})\leq k$ for any $i,j=0,\cdots,n$.
Firstly, define a metric $d_{S_{k}}$ on $P_{k}(N)$ to be the path metric whose restriction to each simplex is the standard spherical metric on the unit sphere by mapping $\sum_{i=0}^{n}t_{i}z_{i}$ to
$\left(\frac{t_{0}}{\sqrt{\sum_{i=0}^{n}t^{2}_{i}}},\cdots,\frac{t_{n}}{\sqrt{\sum_{i=0}^{n}t^{2}_{i}}}\right),$
namely,
$d_{S_{k}}\left(\sum_{i=0}^{n}t_{i}z_{i},\sum_{i=0}^{n}t^{\prime}_{i}z_{i}\right)=\frac{2}{\pi}\arccos\left(\frac{\sum_{i=0}^{n}t_{i}t^{\prime}_{i}}{\sqrt{(\sum_{i=0}^{n}t^{2}_{i})(\sum_{i=0}^{n}t^{\prime 2}_{i})}}\right),$
where $t_{i},t^{\prime}_{i}\in[0,1]$ and $\sum_{i=0}^{n}t_{i}=\sum_{i=0}^{n}t^{\prime}_{i}=1$, and the metric $d_{S_{k}}$ between different connected components is defined to be infinity. Unfortunately, the natural include map from $(N,d_{N})$ to $(P_{k}(N),d_{S_{k}})$ is generally not a coarsely equivalent map. For this reason, we now introduce the following modified metric on $P_{k}(N)$ coming from [18, Definition 7.2.8].
Define a metric $d_{P_{k}}$ on $P_{k}(N)$ to be
$d_{P_{k}}(x,y)=\inf\left\{\sum_{i=0}^{n}d_{S_{k}}(x_{i},y_{i})+\sum_{i=0}^{n-1}d_{N}(y_{i},x_{i+1})\right\},$
for all $x,y\in P_{k}(N)$, here the infimum is taken over all sequences of the form $x=x_{0},y_{0},x_{1},y_{1},\cdots,x_{n},y_{n}=y$, where $x_{1},\cdots,x_{n},y_{0},\cdots,y_{n-1}\in N$.
Then by Lemma 7.2.9 and Proposition 7.2.11 in [18], we have the following lemma.
Lemma 3.20.
For any $k\geq 0$, the metric spaces $(N,d_{N})$ and $(P_{k}(N),d_{P_{k}})$ are coarsely equivalent. More precisely, let $i_{k}:N\rightarrow P_{k}(N)$ be the inclusion map and define $j_{k}:P_{k}(N)\rightarrow N$ by mapping $\sum_{i=0}^{n}t_{i}z_{i}$ to $z_{i}$ for some $t_{i}\neq 0$, then $i_{k}$ and $j_{k}$ are coarse maps and
$\max\{d_{N}(j_{k}i_{k}(z),z),d_{P_{k}}(i_{k}j_{k}(x),x)\}\leq 1,$
for any $z\in N$, $x\in P_{k}(N)$.
For $k_{1}\leq k_{2}$, the natural inclusion map
$i_{k_{1}k_{2}}:P_{k_{1}}(N)\rightarrow P_{k_{2}}(N)$
is a contractible coarse map. Then they induce the following homomorphisms by Lemma 3.6 and Lemma 3.13:
$ad_{i_{k_{1}k_{2}},\ast}:K_{\ast}(C_{f}^{\ast}(P_{k_{1}}(N),A))\rightarrow K_{\ast}(C_{f}^{\ast}(P_{k_{2}}(N),A)),$
$Ad_{i_{k_{1}k_{2}},\ast}:K_{\ast}(C_{L,f}^{\ast}(P_{k_{1}}(N),A))\rightarrow K_{\ast}(C_{L,f}^{\ast}(P_{k_{2}}(N),A)).$
Thus we obtain two inductive systems of abelian groups:
$\{K_{\ast}(C_{f}^{\ast}(P_{k}(N),A)),ad_{kk^{\prime},\ast}\}\>\>\text{and}\>\>\{K_{\ast}(C_{L,f}^{\ast}(P_{k}(N),A)),Ad_{kk^{\prime},\ast}\}.$
Consider the evaluation at zero map:
$e:C_{L,f}^{\ast}(P_{k}(N),A)\rightarrow C_{f}^{\ast}(P_{k}(N),A),\>u\mapsto u(0)$
which induces a homomorphism on $K$-theory level
$e_{\ast}:K_{\ast}(C_{L,f}^{\ast}(P_{k}(N),A))\rightarrow K_{\ast}(C_{f}^{\ast}(P_{k}(N),A)),$
satisfying $ad_{i_{k_{1}k_{2}},\ast}\circ e_{\ast}=e_{\ast}\circ Ad_{i_{k_{1}k_{2}},\ast}$.
Definition 3.21.
Let $X$ be a metric space, a net in $X$ is a subset $N_{X}$ satisfying that there exists $C\geq 1$ such that
  1. (1)
    $d(z,z^{\prime})\geq C$ for any $z,z^{\prime}\in N_{X}$;
  2. (2)
    for any $x\in X$, there exists $z\in N_{X}$ such that $d(z,x)\leq C$.
Remark 3.22.
By Zorn’s lemma, any proper metric space admits a locally finite net. And any net $N_{X}$ is coarsely equivalent to $X$.
Now, we are ready to show the coarse Baum-Connes conjecture with filtered coefficients in a filtered $C^{\ast}$-algebra $A$ as follow:
Conjecture 3.23.
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphisms
$e_{\ast}:\lim_{k\rightarrow\infty}K_{\ast}(C_{L,f}^{\ast}(P_{k}(N_{X}),A))\rightarrow\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X}),A))$
are isomorphic between abelian groups.
Remark 3.24.
We have the following comments for the above conjecture.
  1. (1)
    The conjecture is independent of the choice of locally finite nets $N_{X}$ of $X$. More generally, the coarse Baum-Connes conjecture with filtered coefficients is invariant under coarse equivalences by Corollary 3.8.
  2. (2)
    By Corollary 3.8 and Lemma 3.20, we have that
    $\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X}),A))\cong K_{\ast}(C_{f}^{\ast}(X,A)).$
    Thus the coarse Baum-Connes conjecture with filtered coefficients provides a formula to compute the $K$-theory of Roe algebras with filtered coefficients.
  3. (3)
    If $A$ is equipped with the trivial filtration, namely, $A_{r}=A$ for all $r>0$, then the conjecture with filtered coefficients in $A$ is also called the coarse Baum-Connes conjecture with coefficients in $A$. In particular, the coarse Baum-Connes conjecture with filtered coefficients in $\mathbb{C}$ is so called the coarse Baum-Connes conjecture (cf. [8], [19]).
At the last of this section, we consider a uniform version of the coarse Baum-Connes conjecture with filtered coefficients.
Let $(X_{i},d_{i})_{i\in\mathbb{N}}$ be a sequence of metric spaces. Define the separated coarse union of $(X_{i})_{i\in\mathbb{N}}$ to be a metric space $(\sqcup_{\mathbb{N}}X_{i},d)$ by
(3.3) $d(x,x^{\prime})=\left\{\begin{array}[]{lcl}d_{i}(x,x^{\prime}),&\mbox{for}&x,x^{\prime}\in X_{i}\\ \infty,&\mbox{for}&x\in X_{i},x^{\prime}\in X_{i^{\prime}},i\neq i^{\prime}.\end{array}\right.$
Then $C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure in $\prod_{i\in\mathcal{I}}C_{f}^{\ast}(X_{i},A)$ of the set
$\bigcup_{r>0}\left(\prod_{i\in\mathbb{N}}C_{f}^{\ast}(X_{i},A)_{r}\right).$
And the localization algebra $C_{L,f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ is the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(\sqcup_{\mathbb{N}}X_{i},A)$ satisfying that
$\lim_{t\rightarrow\infty}\sup_{i\in\mathbb{N}}(\operatorname{prop}(u_{i}(t)))=0.$
where $u_{i}(t)$ is the $i$th coordinate of $u(t)$ in $C_{f}^{\ast}(X_{i},A)$.
We call the following conjecture as the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$.
Conjecture 3.25.
Let $X$ be a proper metric space and $N_{X}$ be a locally finite net in $X$, then the following homomorphisms
$e_{\ast}:\lim_{k\rightarrow\infty}K_{\ast}(C_{L,f}^{\ast}(\sqcup_{\mathbb{N}}\left(P_{k}(N_{X})\right),A))\rightarrow\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(\sqcup_{\mathbb{N}}\left(P_{k}(N_{X})\right),A))$
are isomorphic between abelian groups, namely, the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $\sqcup_{\mathbb{N}}X$.
Remark 3.26.
Actually, the uniform coarse Baum-Connes conjecture with coefficients in $A$ equipped with the trivial filtration is equivalent to the quantitative coarse Baum-Connes conjecture with coefficients in $A$, please refer to [23] for more details.
4. Product metric spaces
In this section, we will prove the coarse Baum-Connes conjecture with filtered coefficients for the product metric space of two metric spaces providing each of them satisfies the conjecture with filtered coefficients.
4.1. The coarse Baum-Connes conjecture along one direction with filtered coefficients
In this subsection, we build a bridge called the localization algebra along one direction with filtered coefficients connecting the coarse Baum-Connes conjecture with filtered coefficients for products and the conjecture with filtered coefficients for one component.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two proper metric spaces, then the product space $X\times Y$ equipped with the following metric
$d((x,y),(x^{\prime},y^{\prime}))=\max\{d_{X}(x,x^{\prime}),d_{Y}(y,y^{\prime})\}\>\>\textrm{for}\>\>x,x^{\prime}\in X\>\>\textrm{and}\>\>y,y^{\prime}\in Y,$
is also a proper metric space.
Remark 4.1.
If $d^{\prime}$ is another metric defined on $X\times Y$ that satisfies
  • $d^{\prime}((x,y),(x^{\prime},y))=d_{X}(x,x^{\prime})$ and $d^{\prime}((x,y),(x,y^{\prime}))=d_{Y}(y,y^{\prime})$ for any $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$;
  • $d^{\prime}((x,y),(x^{\prime},y^{\prime}))\geq\max\{d_{X}(x,x^{\prime}),d_{Y}(y,y^{\prime})\}$ for any $x,x^{\prime}\in X$ and $y,y^{\prime}\in Y$.
Then $d^{\prime}$ is bi-Lipschitz equivalent to $d$ as defined above, namely,
$d((x,y),(x^{\prime},y^{\prime}))\leq d^{\prime}((x,y),(x^{\prime},y^{\prime}))\leq 2d((x,y),(x^{\prime},y^{\prime})).$
For examples, we can define $d^{\prime}((x,y),(x^{\prime},y^{\prime}))=(d_{X}(x,x^{\prime})^{p}+d_{Y}(y,y^{\prime})^{p})^{1/p}$ for any $p\geq 1$.
Definition 4.2.
Let $(X,d_{X})$ and $(Y,d_{Y})$ be two proper metric spaces with two countable dense subsets $Z_{X}$ and $Z_{Y}$, respectively. Let $T$ be an operator acting on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H$. The propagation along $X$ of $T$, denoted by $\operatorname{prop}_{X}(T)$, is defined to be
$\sup\{d_{X}(x,x^{\prime}):\>\text{there exist $y,y^{\prime}\in Y$ such that $\left(\left(x,y\right)\left(x^{\prime},y^{\prime}\right)\right)\in\operatorname{supp}(T)$}\}.$
Definition 4.3.
Let $X$ and $Y$ be as above, the localization algebra along $X$ with filtered coefficients in $A$ of $X\times Y$, denoted by $C^{\ast}_{L,X,f}(X\times Y,A)$, is defined to be the norm closure of the $\ast$-algebra consisting of all bounded and uniformly continuous functions $u:[0,\infty)\rightarrow C_{f}^{\ast}(X\times Y,A)$ satisfying that
$\lim_{t\rightarrow\infty}\operatorname{prop}_{X}\left(u\left(t\right)\right)=0.$
Remark 4.4.
Similar definitions as above are appeared in [1][3].
Let $X$, $X^{\prime}$ and $Y$ be three proper metric spaces. For a uniformly continuous coarse map $g:X\rightarrow X^{\prime}$, then $g\times Id:X\times Y\rightarrow X^{\prime}\times Y$ is also a uniformly continuous coarse map. Thus by Lemma 3.13, there exists a $\ast$-homomorphism $Ad_{g\times Id}:C^{\ast}_{L,f}(X\times Y,A)\rightarrow C^{\ast}_{L,f}(X^{\prime}\times Y,A)$ defined by $V_{g\times Id}$. Moreover, for $u\in C^{\ast}_{L,X,f}(X\times Y,A)$, we have that
$\operatorname{prop}_{X^{\prime}}(V_{g\times Id}(t)u(t)V^{\ast}_{g\times Id}(t))\leq w_{g}(\operatorname{prop}_{X}(u(t)))+2\delta_{k},$
for any $t\in[k,k+1]$, where $\lim_{k\rightarrow\infty}\delta_{k}=0$. Thus $Ad_{g\times Id}$ can be extended to a $\ast$-homomorphism from $C^{\ast}_{L,X,f}(X\times Y,A)$ to $C^{\ast}_{L,X^{\prime},f}(X^{\prime}\times Y,A)$.
By the similar proofs of Lemma 3.17 and Lemma 3.19, we also have the following strongly Lipschitz homotopy invariance and Mayer-Vietoris six-term exact sequence for the $K$-theory of localization algebra along $X$ with filtered coefficients.
Lemma 4.5.
Let $X$, $X^{\prime}$ and $Y$ be three proper metric spaces. If $X$ is strongly Lipschitz homotopy equivalent to $X^{\prime}$, then $K_{\ast}(C^{\ast}_{L,X,f}(X\times Y,A))$ is naturally isomorphic to $K_{\ast}(C^{\ast}_{L,{X^{\prime}},f}(X^{\prime}\times Y,A))$.
Lemma 4.6.
Let $X=X_{1}\cup X_{2}$ be a uniformly excisive cover. If there exists $\varepsilon_{0}>0$ such that the $\varepsilon_{0}$-neighborhoods of $X_{1},X_{2},X_{1}\cap X_{2}$ are strongly Lipschitz homotopy equivalent to $X_{1},X_{2},X_{1}\cap X_{2}$, respectively, then there is a Mayer-Vietoris six-term exact sequence
${K_{0}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{0}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{0}(C^{\ast}_{L,{X},f}((X)\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,{X},f}((X)\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{i=1,2}K_{1}(C^{\ast}_{L,{X_{i}},f}(X_{i}\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K_{1}(C^{\ast}_{L,X_{1,2},f}(X_{1,2}\times Y,A)),\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where $X_{1,2}=X_{1}\cap X_{2}$.
Now we introduce the coarse Baum-Connes conjecture along $X$ with filtered coefficients for $X\times Y$.
Conjecture 4.7.
Let $X$, $Y$ be two proper metric spaces and $N_{X}$ be a locally finite net in $X$. Then the evaluation at zero map $e$ induces the following isomorphisms
$e_{\ast}:\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))\rightarrow\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times Y,A)).$
Lemma 4.8.
If two proper metric spaces $Y$ and $Y^{\prime}$ are coarsely equivalent, then the coarse Baum-Connes conjecture along $X$ with filtered coefficients holds for $X\times Y$ if and only if it holds for $X\times Y^{\prime}$.
Proof.
Let $g:Y\rightarrow Y^{\prime}$ be a coarsely equivalent map, then $Id\times g:P_{k}(N_{X})\times Y\rightarrow P_{k}(N_{X})\times Y^{\prime}$ is also coarsely equivalent map which is uniformly continuous on $P_{k}(N_{X})$. Then by Corollary 3.8, we can show that $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y,A))$ are naturally isomorphic to $K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y^{\prime},A))$ and $K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X})\times Y^{\prime},A))$, respectively, for any $k\in\mathbb{N}$. Thus we completed the proof. ∎
Recall from Lemma 3.9 that $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration induced by the propagation on $Y$ and the filtration on $A$. We have the following relation between the coarse Baum-Connes conjecture along $X$ with filtered coefficients and the coarse Baum-Connes conjecture with filtered coefficients.
Proposition 4.9.
Let $X$, $Y$ be two proper metric spaces and $A$ be a filtered $C^{\ast}$-algebra acting on a Hilbert space $H_{A}$. If the coarse Baum-Connes conjecture with filtered coefficients in $C_{f}^{\ast}(Y,A)$ holds for $X$, then the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ holds for $X\times Y$.
Proof.
Let $H$ be a separable Hilbert space and $U:H\rightarrow H\otimes H$ be a unitary. Fixed a unit vector $h_{0}\in H$ and let $P:H\rightarrow H,h\mapsto\langle h,h_{0}\rangle h_{0}$ which is a rank-one projection. The algebra $C_{f}^{\ast}(X\times Y,A)$ acts on $\ell^{2}(Z_{X})\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$ and the algebra $C^{\ast}_{f}(X,C^{\ast}(Y,A))$ acts on $\ell^{2}(Z_{X})\otimes H\otimes\ell^{2}(Z_{Y})\otimes H\otimes H_{A}$, where $Z_{X}$ and $Z_{Y}$ are two countable dense subsets in $X$ and $Y$, respectively. Define two $\ast$-homomorphisms:
$\begin{split}&\phi:C^{\ast}_{f}(X,C^{\ast}(Y,A))\rightarrow C_{f}^{\ast}(X\times Y,A),T_{(x,y),(x^{\prime},y^{\prime})}\mapsto(U^{\ast}\otimes I_{H_{A}})T_{(x,y),(x^{\prime},y^{\prime})}(U\otimes I_{H_{A}});\\ &\psi:C_{f}^{\ast}(X\times Y,A)\rightarrow C^{\ast}_{f}(X,C^{\ast}(Y,A)),S_{(x,y),(x^{\prime},y^{\prime})}\mapsto S_{(x,y),(x^{\prime}y^{\prime})}\otimes P.\end{split}$
Then the homomorphism $\phi\circ\psi:C_{f}^{\ast}(X\times Y,A)\rightarrow C_{f}^{\ast}(X\times Y,A)$ satisfies that
$\phi\circ\psi(S)_{(x,y),(x^{\prime},y^{\prime})}=(U^{\ast}V\otimes I_{H_{A}})S_{(x,y),(x^{\prime},y^{\prime})}(V^{\ast}U\otimes I_{H_{A}}),$
where $V:H\rightarrow H\otimes H$ maps $h$ to $h\otimes h_{0}$. Since the operator $U^{\ast}V\otimes I_{H_{A}}$ is an isometry in the multiplier algebra of $C^{\ast}_{f}(X\times Y,A)$. Thus, by Lemma 2.1, $\phi_{\ast}\circ\psi_{\ast}$ is the identity homomorphism on $K_{\ast}(C_{f}^{\ast}(X\times Y,A))$. Similarly, we can define the above two homomorphisms for localization algebras. Therefore, we have the following commutative diagram:
${\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\ast}}$$\scriptstyle{e_{\ast}}$${\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\psi_{\ast}}$${\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X}),C^{\ast}(Y,A)))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\ast}}$$\scriptstyle{e_{\ast}}$${\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{f}(P_{k}(N_{X}),C^{\ast}(Y,A)))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\ast}}$${\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times Y,A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{e_{\ast}}$${\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times Y,A)),}$
where $N_{X}$ is a locally finite net in $X$. Since the compositions of two vertical homomorphisms are all the identity homomorphism, thus if the middle horizontal map is an isomorphism, then the top horizontal map is injective and the bottom horizontal map is surjective which completes the proof. ∎
Remark 4.10.
For two proper metric spaces $X$ and $Y$, the algebra $C^{\ast}_{f}(X\times Y,A)$ is generally not isomorphic to the algebra $C^{\ast}_{f}(X,C^{\ast}(Y,A))$. For example, If both $X$ and $Y$ are the separated coarse union of a sequence of single points and $A=\mathbb{C}$, then $C^{\ast}_{f}(X\times Y,A)=\ell^{\infty}(X,\ell^{\infty}(Y,\mathfrak{K}(H)))$ and $C^{\ast}_{f}(X,C^{\ast}(Y,A))=\ell^{\infty}(X,\mathfrak{K}(H)\otimes\ell^{\infty}(Y,\mathfrak{K}(H)))$ that are not isomorphic.
Lemma 4.11.
Let $X$ be a proper metric space, $\sqcup_{\mathbb{N}}X$ be the separated coarse union of $X$. Then $X$ satisfies the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ if and only if $X\times\square$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$, where $\square$ is the separated coarse union of a sequence of single points.
Proof.
Firstly, we have that $\sqcup_{\mathbb{N}}X=X\times\square$, thus
$C_{f}^{\ast}(\sqcup_{\mathbb{N}}X,A)=C_{f}^{\ast}(X\times\square,A);\>\>C_{L,f}^{\ast}(\sqcup_{\mathbb{N}}X,A)=C_{L,f}^{\ast}(X\times\square,A).$
Secondly, for an operator $T\in C_{f}^{\ast}(X\times\square,A)$, we have that $\operatorname{prop}_{X}(T)=\operatorname{prop}(T)$, thus
$C^{\ast}_{L,f}(X\times\square,A)=C^{\ast}_{L,X,f}(X\times\square,A).$
Finally, for a locally finite net $N_{X}$ in $X$, we have that $P_{k}(\sqcup_{\mathbb{N}}X)=\sqcup_{\mathbb{N}}P_{k}(X)$ for any non-negative integer $k$. Thus we have that
$C^{\ast}_{L,f}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C^{\ast}_{L,P_{k}(X),f}(P_{k}(X)\times\square,A);\>\>C_{f}^{\ast}(P_{k}(\sqcup_{\mathbb{N}}X),A)=C_{f}^{\ast}(P_{k}(X)\times\square,A),$
which completes the proof. ∎
Combing the above lemma with Lemma 4.9, we have the following corollary.
Corollary 4.12.
Let $X$ be a proper metric space. If the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $X$, then the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X$.
Remark 4.13.
In [23], the authors introduced the quantitative coarse Baum-Connes conjecture with coefficients which refines the original conjecture and proved this conjecture is equivalent to the uniform coarse Baum-Connes conjecture with coefficients, thus by the above corollary, the coarse Baum-Connes conjecture with filtered coefficients implies that conjecture.
4.2. Main theorems
Definition 4.14.
A locally finite metric space $N$ is called to have bounded geometry if for every $R>0$, $\sup_{x\in N}\#B(x,R)$ is finite, where #$\#B(x,R)$ is the cardinality of the $R$-ball centered at $x$. A proper metric space $X$ is called to have bounded geometry if there exists a net $N_{X}$ with bounded geometry.
Firstly, let us discuss the Rips complexes of product metric spaces.
Lemma 4.15.
Let $N_{1}$ and $N_{2}$ be two locally finite metric spaces with bounded geometry, then $P_{k}(N_{1}\times N_{2})$ is strongly Lipschitz homotopy equivalent to and coarsely equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$ for any $k\in\mathbb{N}$.
Proof.
Since $N_{1}$ and $N_{2}$ have bounded geometry, thus the dimension of simplicial complexes $P_{k}(N_{1})$ and $P_{k}(N_{2})$ is less than some positive integer $m_{k}$. Define two maps:
$\rho:P_{k}\left(N_{1}\times N_{2}\right)\rightarrow P_{k}\left(N_{1}\right)\times P_{k}\left(N_{2}\right)$
by
$\sum_{i,j}t_{ij}\left(x_{i},y_{j}\right)\mapsto\left(\sum_{i}\left(\sum_{j}t_{ij}\right)x_{i},\sum_{j}\left(\sum_{i}t_{ij}\right)y_{j}\right),$
and
$\varphi:P_{k}(N_{1})\times P_{k}(N_{2})\rightarrow P_{k}(N_{1}\times N_{2})$
by
$\left(\sum_{i}t_{i}x_{i},\sum_{j}s_{j}y_{j}\right)\mapsto\sum_{i,j}t_{i}s_{j}\left(x_{i},y_{j}\right),$
where $x_{i}\in N_{1}$ and $y_{j}\in N_{2}$. Then we have that
$d_{P_{k}}(\rho(x),\rho(y))\leq 4m_{k}^{2}d_{P_{k}}(x,y);\>\>d_{P_{k}}(\varphi(z),\varphi(w))\leq 2d_{P_{k}}(z,w)$
for any $x,y\in P_{k}(N_{1}\times N_{2})$ and $z,w\in P_{k}(N_{1})\times P_{k}(N_{2})$ by direct computations which imply that $\rho$ and $\varphi$ are two coarse and Lipschitz maps. Moreover, $\rho\circ\varphi=id$ and $\varphi\circ\rho$ is strongly Lipschitz homotopic to the identity map by the linear combination of them. Thus, $P_{k}(N_{1}\times N_{2})$ is strongly Lipschitz homotopy equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. And we have that $d(\rho\circ\varphi(z),z)=0$ and $d(\varphi\circ\rho(x),x)\leq 1$, thus $P_{k}(N_{1}\times N_{2})$ is coarsely equivalent to $P_{k}(N_{1})\times P_{k}(N_{2})$. ∎
Corollary 4.16.
Let $X$, $Y$ be two proper metric spaces with bounded geometry and $A$ be a filtered $C^{\ast}$-algebra. Then $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$ if and only if the following map
$\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{k}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}(P_{k}(N_{X})\times P_{k}(N_{Y}),A)).$
is an isomorphism, where $N_{X}$ and $N_{Y}$ are two nets with bounded geometry in $X$ and $Y$, respectively.
Proof.
For any $k^{\prime}\geq k$, we have the following commutative diagram:
[…]
where $\varphi$ is defined in the proof of the above lemma.
Then we further have the following commutative diagram:
[…]
Thus, the proof is completed by combining Corollary 3.8, Lemma 3.17 and the above lemma. ∎
Secondly, we give the following lemma which plays a crucial role in the proof of main theorems.
Let $X$, $Y$ be two proper metric spaces and $A$ be a filtered $C^{\ast}$-algebra. Then we have the following inclusion map:
$\tau:C^{\ast}_{L,f}(X\times Y,A)\rightarrow C^{\ast}_{L,X,f}(X\times Y,A).$
Lemma 4.17.
Assume $X$ has bounded geometry. Let $N_{X}$ be a net with bounded geometry in $X$ and $N_{Y}$ be a locally finite net in $Y$. If $Y$ satisfies the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$, then for each $k>0$
$\tau_{\ast}:\lim_{d\rightarrow\infty}K_{\ast}\left(C^{\ast}_{L,f}\left(P_{k}\left(N_{X}\right)\times P_{d}\left(N_{Y}\right),A\right)\right)\rightarrow\lim_{d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))$
is an isomorphism.
Proof.
Since $N_{X}$ has bounded geometry, thus $P_{k}(N_{X})$ is a finite-dimensional simplicial complex for each $k$. Let $P_{k}(N_{X})^{(n)}$ be its $n$-dimensional skeleton. We will prove the lemma for $P_{k}(N_{X})^{(n)}$ by induction on $n$.
When $n=0$, we have the following commutative diagram:
${\lim_{d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(\sqcup_{\mathbb{N}}P_{d}(N_{Y}),A))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\tau_{\ast}\>\>\>\>\>\>\>\>\>\>\>\>\>\>\>}$$\scriptstyle{e_{\ast}}$${\lim_{d\rightarrow\infty}K_{\ast}(C_{ub}([0,\infty),C_{f}^{\ast}(\sqcup_{\mathbb{N}}P_{d}(N_{Y}),A)))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{e_{\ast}}$${\lim_{d\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(\sqcup_{\mathbb{N}}P_{d}(N_{Y}),A)),}$
where $\sqcup_{\mathbb{N}}P_{d}(N_{Y})$ is the separated coarse union. Then by Lemma 2.3 and Lemma 3.10, we obtain that the right-hand side vertical map $e_{\ast}$ is an isomorphism. Thus by the assumption, $\tau_{\ast}$ is an isomorphism for the case of $n=0$.
Assume by induction that the lemma holds for $n=l-1$. Next we will prove the lemma also holds for $n=l$. For each simplex $\triangle$ of dimension $l$ in $P_{k}(N_{X})$, let $c(\triangle)$ be the center of $\triangle$. Then define
$\triangle_{1}=\{x\in\triangle:d(x,c(\triangle))\leq 1/10\},\>\>\triangle_{1}=\{x\in\triangle:d(x,c(\triangle))\geq 1/10\}.$
And let
$X_{1}=\cup\{\triangle_{1}:\text{$\triangle$ is a simplex of dimension $l$}\};$
$X_{2}=\cup\{\triangle_{2}:\text{$\triangle$ is a simplex of dimension $l$}\},$
then $P_{k}(N_{X})^{(l)}=X_{1}\cup X_{2}$ and $X_{1}\cap X_{2}$ is strongly Lipschitz homotopy equivalent to the disjoint union of the boundaries of all $l$-dimensional simplexes in $P_{k}(N_{X})$. Besides, $X_{1}$ and $X_{2}$ are strongly Lipschitz homotopy equivalent to {c(): $\{c(\triangle):\text{$\triangle$ is a simplex of dimension $l$}\}$ and $P_{k}(N_{X})^{(l-1)}$, respectively. Thus by Lemma 3.17, Lemma 3.19, Lemma 4.5, Lemma 4.6 and the five lemma, we completed the proof. ∎
Now, let us state and prove our main theorems.
Theorem 4.18.
Let $X$, $Y$ be two proper metric spaces with bounded geometry and $A$ be a filtered $C^{\ast}$-algebra. If the following two conditions are satisfied:
  1. (1)
    the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ holds for $X\times Y$;
  2. (2)
    the uniform coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $Y$.
Then the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X\times Y$.
Proof.
Let $N_{X}$, $N_{Y}$ be two nets with bounded geometry in $X$ and $Y$, respectively. Then by Condition 1 and Lemma 3.20, Lemma 4.8, we have that $X\times P_{d}(N_{Y})$ satisfies the coarse Baum-Connes conjecture along $X$ with filtered coefficients in $A$ for any $d\in\mathbb{N}$, namely, the following map is an isomorphism
$\lim_{k\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times P_{d}(N_{Y}),A)).$
On the other hand, by Condition 2 and Lemma 4.17, we have the following isomorphism for any $k\in\mathbb{N}$:
$\lim_{d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))\stackrel{{\scriptstyle\tau_{\ast}}}{{\longrightarrow}}\lim_{d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,P_{k}(N_{X}),f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A)).$
Combining the above two isomorphisms, we obtain an isomorphism:
$\lim_{k,d\rightarrow\infty}K_{\ast}(C^{\ast}_{L,f}(P_{k}(N_{X})\times P_{d}(N_{Y}),A))\stackrel{{\scriptstyle e_{\ast}}}{{\longrightarrow}}\lim_{k,d\rightarrow\infty}K_{\ast}(C_{f}^{\ast}(P_{k}(N_{X})\times P_{d}(N_{Y}),A)).$
which implies that $X\times Y$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$ by Corollary 4.16. ∎
Recall that $C^{\ast}_{f}(Y,A)$ is a filtered $C^{\ast}$-algebra equipped with a filtration defined in Lemma 3.9 and $\square$ is the separated coarse union of a sequence of single points. Combining the above theorem with Proposition 4.9 and Corollary 4.12, we have the following theorem which implies that the coarse Baum-Connes conjecture with filtered coefficients is closed under products.
Theorem 4.19.
Let $X$, $Y$ be two proper metric spaces with bounded geometry and $A$ be a filtered $C^{\ast}$-algebra. If the following two conditions are satisfied:
  1. (1)
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(Y,A)$ holds for $X$;
  2. (2)
    the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}_{f}(\square,A)$ holds for $Y$.
Then the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X\times Y$.
In particular, if the coarse Baum-Connes conjecture with filtered coefficients holds for $X$ and $Y$, then the coarse Baum-Connes conjecture with filtered coefficients holds for $X\times Y$.
When $A=\mathbb{C}$, we have that $C^{\ast}_{f}(\square,\mathbb{C})=\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$ equipped with the trivial filtration, namely, $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))_{r}=\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$ for all $r>0$. Thus, we obtain the following corollary which enlarges the class of product metric spaces that satisfy the coarse Baum-Connes conjecture.
Corollary 4.20.
Let $X$ and $Y$ be two proper metric spaces with bounded geometry. If the followings are satisfied:
  1. (1)
    $X$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $C^{\ast}(Y)$;
  2. (2)
    $Y$ satisfies the coarse Baum-Connes conjecture with coefficients in $\ell^{\infty}(\mathbb{N},\mathfrak{K}(H))$.
Then the coarse Baum-Connes conjecture holds for $X\times Y$.
In particular, if $X$ and $Y$ satisfy the coarse Baum-Connes conjecture with filtered coefficients, then $X\times Y$ satisfies the coarse Baum-Connes conjecture.
5. Applications
In this section, we state that three ways mentioned in Introduction to the coarse Baum-Connes conjecture also work for the conjecture with filtered coefficients, thus, their product metric spaces satisfy the conjecture with filtered coefficients by Theorem 4.19. The reason is that all morphisms appeared in the proofs of those three ways are the identity map on filtered coefficients. The results in this section recover Fukaya and Oguni’s result in [5] and also contain some new examples satisfying the coarse Baum-Connes conjecture.
5.1. Metric spaces with finite asymptotic dimension
Definition 5.1.
Call a metric space $X$ has finite asymptotic dimension, if there is a non-negative integer $m<\infty$ satisfying that for any $r>0$, there exists a uniformly bounded cover $\{U_{i}\}_{i\in I}$ of $X$ such that no ball with radius $r$ in $X$ intersects more than $m+1$ members of $\{U_{i}\}_{i\in I}$.
Remark 5.2.
Every polycyclic group (which is solvable) equipped with a proper metric has finite asymptotic dimension (cf. [13, Chapter 2]).
In [21], Yu invented the quantitative $K$-theory to prove that the coarse Baum-Connes conjecture holds for metric spaces with finite asymptotic dimension.
Let $X$ be a proper metric space and $A$ be a filtered $C^{\ast}$-algebra, then the Roe algebra $C_{f}^{\ast}(X,A)$ is also a filtered $C^{\ast}$-algebra equipped with a filtration defined by
$C_{f}^{\ast}(X,A)_{s}=\{T\in C_{f}^{\ast}(X,A):\operatorname{prop}(T)\leq s\},$
for $s\geq 0$ (please compare it with Lemma 3.9). Then by using the quantitative $K$-theory to $C_{f}^{\ast}(X,A)$ with the above filtration, we can obtain the following result.
Proposition 5.3.
Let $X$ be a proper metric space with finite asymptotic dimension and $A$ be a filtered $C^{\ast}$-algebra, then the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X$.
5.2. Metric spaces with coarse embeddability into Hilbert space and CE-by-CE extensions
Definition 5.4.
A proper metric space $X$ is called to be coarsely embedded into Hilbert space $H$, if there exists a map $f:X\rightarrow H$ and two non-deceasing functions $\rho_{\pm}:[0,\infty)\rightarrow[0,\infty)$ with $\lim_{t\rightarrow\infty}\rho_{-}(t)=\infty$ such that
$\rho_{-}(d(x,x^{\prime}))\leq\parallel f(x)-f(x^{\prime})\parallel\leq\rho_{+}(d(x,x^{\prime})),$
for all $x,x^{\prime}\in X$.
In [22], Yu used the Dirac-dual-Dirac method to prove that the coarse Baum-Connes conjecture holds for any metric space which admits a coarse embedding into Hilbert space (see [18, Chapter 12] for a simplified proof). We can take tensor products of the Dirac map and the dual-Dirac map with the identity map on filtered coefficients in the Dirac-dual-Dirac method, then we obtain the following result.
Proposition 5.5.
Let $X$ be a proper metric space with bounded geometry and $A$ be a filtered $C^{\ast}$-algebra. If $X$ can be coarsely embedded into Hilbert space, then the coarse Baum-Connes conjecture with filtered coefficients in $A$ holds for $X$.
The following notion was introduced by Deng, Wang and Yu in [2] in order to consider the coarse Baum-Connes conjecture for relative expanders. Recall that the coarse union of a sequence of metric spaces $(X_{i})_{i\in\mathbb{N}}$ is a disjoint union $\sqcup_{i\in\mathbb{N}}X_{i}$ equipped with a metric $d$ that restricts to the original metric on each $X_{i}$ and satisfies $d(X_{n},(\sqcup_{\mathbb{N}}X_{i})\setminus X_{n})\rightarrow\infty$ as $n\rightarrow\infty$.
Definition 5.6.
A sequence of group extensions $(1\rightarrow N_{m}\rightarrow\Gamma_{m}\rightarrow Q_{m}\rightarrow 1)_{m\in\mathbb{N}}$ is called to be a CE-by-CE extension, if both the coarse union of $(N_{m})_{m\in\mathbb{N}}$ and the coarse union of $(Q_{m})_{m\in\mathbb{N}}$ admit a coarse embedding into Hilbert space.
Using the same arguments (relies on the Dirac-dual -Dirac method) in [2], we can obtain the following lemma.
Proposition 5.7.
Let $(1\rightarrow N_{m}\rightarrow\Gamma_{m}\rightarrow Q_{m}\rightarrow 1)_{m\in\mathbb{N}}$ be a CE-by-CE extension with all coarse unions of $(N_{m})_{m\in\mathbb{N}}$, $(\Gamma_{m})_{m\in\mathbb{N}}$ and $(Q_{m})_{m\in\mathbb{N}}$ have bounded geometry. Then the coarse Baum-Connes conjecture with filtered coefficients in any filtered $C^{\ast}$-algebra holds for the coarse union of $(\Gamma_{m})_{m\in\mathbb{N}}$.
Remark 5.8.
Recently, Deng and Guo in [1] introduced a twisted coarse Baum-Connes conjecture and showed that the coarse Baum-Connes conjecture holds for the coarse union of $(\Gamma_{m})_{m\in\mathbb{N}}$ if coarse unions of $(N_{m})_{m\in\mathbb{N}}$ and $(Q_{m})_{m\in\mathbb{N}}$ satisfy the twisted coarse Baum-Connes conjecture with coefficients which generalizes the main result in [2]. Actually, we could obtain the similar result for the coarse Baum-Connes conjecture with filtered coefficients by considering the uniform version of the conjecture with filtered coefficients.
5.3. Open cones, coarse homotopy and relatively hyperbolic groups
Definition 5.9.
Let $(M,d_{M})$ be a compact metric space with diameter at most two. The open cone over $M$, denoted by $\mathcal{O}M$, is a metric space defined to be the quotient space $(\mathbb{R}_{\geq 0}\times M)/(\{0\}\times M)$ with the following metric
$d((t,x),(s,y))=\lvert t-s\rvert+\min\{t,s\}d_{M}(x,y),$
for any $t,s\in\mathbb{R}_{+}$ and $x,y\in M$.
Definition 5.10.
[6, Definition 2.1] Let $f,g:X\rightarrow Y$ be two coarse maps between proper metric spaces. $f$ and $g$ are called to be coarsely homotopic, if there exists a metric subspace $Z=\{(x,t):0\leq t\leq t_{x}\}$ of $X\times\mathbb{R}$ and a coarse map $h:Z\rightarrow Y$, such that
  • the map from $X$ to $\mathbb{R}$ given by $x\mapsto t_{x}$ satisfies that for any $R\geq 0$, there exists $S\geq 0$ such that $|t_{x}-t_{x^{\prime}}|\leq S$ for all elements $x,x^{\prime}\in X$ with $d(x,x^{\prime})\leq R$,
  • $h(x,0)=f(x)$, $h(x,t_{x})=g(x)$.
The map $f$ is called a coarse homotopy equivalence map if there exists a coarse map $f^{\prime}:Y\rightarrow X$ such that $f^{\prime}f$ and $ff^{\prime}$ are coarsely homotopic to the identities $id_{X}$ and $id_{Y}$, respectively. Call $X$ and $Y$ be coarsely homotopy equivalent if there exists a coarse homotopy equivalence map from $X$ to $Y$.
Every open cone satisfies the coarse Baum-Connes conjecture and the proof relied on an Eilenberg swindle can be applied to the conjecture with filtered coefficients in any filtered $C^{\ast}$-algebras (cf. [24], [8]). Moreover, the conjecture with filtered coefficients is a coarse homotopy invariant (the proof relies on the Mayer-Vietoris six-term exact sequence, cf. [24], [9, Chapter 12]). Thus we have the following theorem.
Proposition 5.11.
If a proper metric space $X$ is coarsely homotopy equivalent to an open cone, then the coarse Baum-Connes conjecture with filtered coefficients in any filtered $C^{\ast}$-algebra holds for $X$.
Example 5.12.
In [6], Fukaya and Oguni provided a class of metric spaces which are coarsely homotopy equivalent to an open cone, including
  1. (1)
    all hyperbolic spaces;
  2. (2)
    Busemann non-positively curved spaces, for example CAT(0)-spaces;
  3. (3)
    systolic groups with the word length metric, especially, Artin groups of almost large type and graphical small cancellation groups;
  4. (4)
    Helly groups with the word length metric, especially, weak Garside groups of finite type and FC-type Artin groups ([11]).
In [4], Fukaya and Oguni proved that if a group is hyperbolic relative to a finite family of infinite subgroups (cf. [4, Definition 2.4]), then this group satisfies the coarse Baum-Connes conjecture provided each subgroup belonging to the family satisfies the conjecture and admits a finite universal space for proper actions. By a similar argument, we can prove the following lemma.
Proposition 5.13.
Let $\Gamma$ is a finitely generated group which is hyperbolic relative to a finite family of infinite subgroups $\{N_{1},\cdots,N_{k}\}$ and $A$ be a filtered $C^{\ast}$-algebra. If every $N_{i}$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$ and admits a finite-dimensional simplicial complex which is a universal space for proper actions, then $\Gamma$ satisfies the coarse Baum-Connes conjecture with filtered coefficients in $A$.
In [5], Fukaya and Oguni showed that the coarse Baum-Connes conjecture holds for products of hyperbolic groups, CAT(00)-groups, polyclic groups and certain relatively hyperbolic groups by consider the coronae of products. But it is generally not known whether the coarse Baum-Connes conjecture is true or not for products when every component satisfies the conjecture.
To sum up this section, by Theorem 4.19, we can obtain that the product spaces of metric spaces appeared in this section satisfy the coarse Baum-Connes conjecture with filtered coefficients, in particular, satisfy the original conjecture. That not only recovers Fukaya and Oguni’s result in [5, Theorem 1] by combining Examples 5.12 with Proposition 5.11 and Proposition 5.13, but also gives some new examples of products satisfying the coarse Baum-Connes conjecture.